### v2

# Revisiting In-Context Learning with Long Context Language Models

Abstract
In-Context Learning (ICL) is a technique by which language models make predictions based on examples provided in their input context. Previously, their context window size imposed a limit on the number of examples that can be shown, making example selection techniques crucial for identifying the maximally effective set of examples. However, the recent advent of Long Context Language Models (LCLMs) has significantly increased the number of examples that can be included in context, raising an important question of whether ICL performance in a many-shot regime is still sensitive to the method of sample selection. To answer this, we revisit these approaches in the context of LCLMs through extensive experiments on 18 datasets spanning 4 tasks. Surprisingly, we observe that sophisticated example selection techniques do not yield significant improvements over a simple random sample selection method. Instead, we find that the advent of LCLMs has fundamentally shifted the challenge of ICL from that of selecting the most effective examples to that of collecting sufficient examples to fill the context window. Specifically, in certain datasets, including all available examples does not fully utilize the context window; however, by augmenting the examples in context with a simple data augmentation approach, we substantially improve ICL performance by 5%.
Revisiting In-Context Learning with Long Context Language Models
Jinheon Baek1thanks: This work was conducted during an internship at Google. For correspondence, contact jinheon.baek@kaist.ac.kr.   Sun Jae Lee2   Prakhar Gupta2 Geunseob (GS) Oh2Siddharth Dalmia2Prateek Kolhar2 KAIST1    Google DeepMind2
1 Introduction
In-Context Learning (ICL) has emerged as a powerful paradigm in natural language processing that enables Language Models (LMs) to learn, adapt, and generalize from examples provided within their input context, eliminating the need for extensive training and parameter updates Brown et al. (2020); Min et al. (2022); von Oswald et al. (2023). However, due to the limited context lengths of earlier LMs (which accommodate only a few thousand tokens), much of previous ICL work has focused on optimizing sample selection strategies Liu et al. (2021); Rubin et al. (2022); Sorensen et al. (2022); An et al. (2023); Mavromatis et al. (2023); Liu et al. (2024). With the advent of Long Context Language Models (LCLMs), which are capable of processing over a million tokens in a single context window, these constraints are significantly relaxed as it enables including a large number of examples to be used in ICL, known as many-shot ICL Agarwal et al. (2024); Bertsch et al. (2024).
Figure 1: Results of various sample selection approaches in many-shot ICL with LCLMs. Approaches include Retrieval that selects examples similar to the target query, Diversity that aims for maximizing example variety, Curriculum that arranges examples in order from easiest to hardest, and Hard that uses only challenging examples, alongside Random that selects examples without any constraints. Results indicate that sample selection methods provide no significant improvement over the naive (random) approach and sometimes perform worse. Meanwhile, Augmentation refers to the approach that generates additional demonstrations and uses them along with original samples for ICL, for low-resource tasks (such as translation, reasoning, and classification) that do not contain enough samples to utilize the full capacity of LCLMs, showing substantial performance gains.
This expansion of context length raises an important question: do previous sample selection strategies, designed for shorter context windows in earlier LMs, generalize to the many-shot ICL regime? To answer this, we systematically revisit existing sample selection strategies by conducting extensive experiments across 18 datasets spanning diverse tasks (namely, classification, translation, summarization, and reasoning) with multiple LCLMs. Our experiments include three types of sample selection methods: relevance, diversity, and difficulty-based sample selection, as outlined in Dong et al. (2023). From these experiments, we uncover novel and surprising findings: contrary to prevailing expectations that carefully selected ICL demonstrations would yield performance improvements, they are similarly effective with a simple random selection approach, offering no statistically meaningful improvements in almost all cases (Figure 1). An additional reason to prefer the naive sample selection approach is that it enables greater efficiency through key-value caching of in-context examples (as the same examples can be reused across multiple queries), unlike sophisticated sample selection methods where the examples vary for each sample.
While the expanded context length in LCLMs allows us to focus less on selecting optimal subsets of examples, it introduces a new challenge: effectively utilizing this expanded capacity when the number of examples is limited. Specifically, in scenarios where available data is sparse (such as low-resource translation or reasoning tasks where annotated data samples are difficult or costly to obtain), the examples available only utilize a small fraction of the full context window. In other words, this mismatch between context capacity and example availability introduces a new direction in ICL research, shifting the focus from optimizing sample selection to maximally utilizing the long context window. To address this, we propose a simple yet effective data augmentation approach to increase the number of in-context examples, which consists of two consecutive steps: (1) generating synthetic examples and (2) filtering out low-quality examples via LCLM prompting prepended with randomly sampled real examples. Then, by adding these augmented data samples to the context, we significantly improve ICL performance.
Moreover, we explore other key factors unique to LCLM-enabled ICL. Specifically, we investigate the capacity of LCLMs to comprehend extremely long context (where a large number of examples up to the context length are present), as well as how they handle scenarios in which some of these examples introduce noise. Through comprehensive analyses, we find that while performance generally improves as the number of in-context examples increases, it eventually plateaus and begins to decline as the context length approaches the limit. This diminishing return highlights the need to carefully balance context length and example quantity within the expanded capacity of LCLMs. In addition, we observe that LCLMs exhibit robustness to noisy examples in relatively simple tasks, but they become vulnerable to noise in more complex scenarios to which they might be less exposed during training, such as extremely low-resource translation tasks.
Overall, we believe our work sheds new light on an important paradigm shift in ICL with LCLMs: the shift from optimizing sample selection to better utilizing extensive context capacity. In particular, our findings suggest that simpler, more efficient random sampling approaches can be as effective as previous sample selection approaches in many-shot settings in most cases, and that data augmentation can significantly improve ICL performance in low-resource tasks. Furthermore, our study paves the way for future research on understanding how to better utilize large context windows and manage the intricacies that arise in extended-context ICL.
2 Examining Sample Selection Methods for In-Context Learning with LCLMs
2.1 Background
We begin with formally introducing LCLMs, followed by describing the setup of ICL with LCLMs.
Long-Context Language Models
A language model (LM), which takes an input sequence of tokens ${\bm{x}}=[x_{1},x_{2},\ldots,x_{n}]$ and generates an output sequence of tokens ${\bm{y}}=[y_{1},y_{2},\ldots,y_{m}]$, can be represented as follows: ${\bm{y}}=\texttt{LM}_{\theta}({\bm{x}})$, where $\theta$ is the set of model parameters that are typically fixed after training due to the high computational costs of fine-tuning. A long-context LM (LCLM) is an advanced LM Reid et al. (2024) that is designed to accommodate sequences with a large number of tokens (e.g., $n$ can exceed 1 million), typically far surpassing the context sizes of earlier LMs.
In-Context Learning with LCLMs
Given a set of $k$ input-output pairs $\{({\bm{x}}_{i},{\bm{y}}_{i})\}_{i=1}^{k}$ as well as an input query ${\bm{x}}^{\prime}$, the goal of ICL is to produce an output ${\bm{y}}=\texttt{LCLM}({\bm{x}}^{\prime}|\{({\bm{x}}_{i},{\bm{y}}_{i})\}_{i=1}^{k})$, where the model (LCLM) uses the contextual examples $\{({\bm{x}}_{i},{\bm{y}}_{i})\}_{i=1}^{k}$ to make predictions for ${\bm{x}}^{\prime}$. In prior research before the advent of LCLMs, the value of $k$ was often limited by the relatively short context lengths of earlier models, which constrained the number of examples that could be utilized for ICL. Subsequently, significant work has focused on developing sample selection techniques to optimize performance within these restricted contexts Liu et al. (2021); Rubin et al. (2022); Sorensen et al. (2022); An et al. (2023); Mavromatis et al. (2023); Liu et al. (2024). In the meantime, the expanded context capacity of LCLMs enables a larger $k$, facilitating many-shot learning with a far greater number of examples.
2.2 Experimental Setup
We now discuss the detailed experimental design.
Tasks and Datasets
We experiment with 18 different datasets across four tasks to evaluate the effectiveness and robustness of various approaches.
  • Translation: This task evaluates the ability of models to translate text from one language to another. We include translations from English to low-resource languages (namely, Bemba, Northern Kurdish, and Ewe) and high-resource languages (Spanish, French, and German) from the FLORES-200 benchmark NLLB et al. (2022), with chrF scores Popovic (2015) as the metric.
  • Summarization: This task assesses the capability of models to generate concise and coherent summaries from articles. We include one widely-used XSum dataset Narayan et al. (2018) and two long-context summarization datasets: ArXiv and GovReport Cohan et al. (2018); Huang et al. (2021). ROUGE-L score is used for evaluation.
  • Reasoning: This task evaluates the capability of models to perform complex reasoning. We use three challenging datasets from Big Bench Hard (BBH) Suzgun et al. (2022) following the experimental setting of Long-Context Frontiers (LOFT) benchmark Lee et al. (2024a).
  • Classification: This task includes challenging benchmark datasets for ICL from Li et al. (2024), particularly designed for classification problems with diverse classes and long inputs.
Figure 2: Detailed results of various sample selection approaches on ICL with LCLMs, such as Gemini Pro (Top), Gemini Flash (Middle), and Llama 3.1 (Bottom), across four different tasks (translation, summarization, reasoning, and extreme classification) with 18 datasets. Each bar represents the averaged performance, with the upper and lower limits indicating standard deviation.
ICL Sample Selection Strategies
To ensure comprehensive coverage of previously explored sample selection strategies, we follow the category of three core dimensions from Dong et al. (2023) (that extensively summarizes around ICL 200 papers). This includes selecting samples based on their diversity, difficulty, and relevance to the query, with the baseline of random sample selection.
  • Naive: This method randomly selects examples from a dataset and uses this initial set of selected examples as ICL demonstrations for all queries.
  • Relevance: This method selects examples that are most similar to the input query to maximize the alignment of ICL demonstrations with the query. To compute semantic similarity between the query and each example, we use an embedding model Lee et al. (2024b).
  • Diversity: This method selects examples that are maximally distinct from each other to capture a broad coverage of features and characteristics within the task space. We first embed each example in a shared embedding space with Lee et al. (2024b) and utilize $k$-means clustering (where $k$ corresponds to the number of desired ICL examples) to group the examples into subcategories. We then select the example closest to each cluster center as the representative to capture a diverse subset of the task features.
  • Difficulty: This method selects examples based on their difficulty. We examine two approaches: the first method (called Curriculum) follows a curriculum learning paradigm where examples are ordered from easiest to hardest; the second one (called Hard) includes only difficult examples, as simpler examples may already be well-understood by models. To assess example difficulty, we use model-based evaluation Liu et al. (2023), which prompts LCLMs 30 times and averages difficulty scores weighted by probabilities.
LCLM Configurations for ICL
We consider LCLMs that support extensive token capacities to evaluate ICL performance in long-context, many-shot ICL scenarios. We focus on models that have context window lengths on the order of millions: Gemini 1.5 Flash, which can process up to 1 million tokens; Gemini 1.5 Pro, which can process up to 2 million tokens Reid et al. (2024). In addition, we also consider the Llama 3.1 70B model Dubey et al. (2024), which, while supporting the comparatively smaller context size of 128K tokens, is still considered an LCLM. For all experiments, we utilize the default hyperparameters for both Gemini and Llama. To provide a comprehensive view of performance under different shots, we vary the number of ICL examples, starting from one and sequentially doubling to 2, 4, 8, 16, 32, and so forth, until reaching either the context size limit or the maximum number of dataset samples, whichever is exhausted first. Furthermore, to ensure the reliability of our results, we conduct multiple runs for each experimental setup: 3 runs for translation and summarization tasks; 10 runs for reasoning and classification tasks. The prompts used to elicit responses from ICL are provided in Appendix A.
2.3 Experimental Results
Table 1: Counting the statistical significance of sophisticated selection approaches over random selection on each experiment instance, by conducting the t-test with 95% confidence threshold. Tran., Summ., Reas, Clas, denote translation, summarization, reasoning, and classification tasks, respectively.
LCLMs Methods Tran. Summ. Reas. Clas. Total
Gemini Pro Relevance 0 / 6 0 / 3 0 / 4 0 / 5 0 / 18
Diversity 0 / 6 0 / 3 1 / 4 2 / 5 3 / 18
Curriculum 1 / 6 0 / 3 0 / 4 1 / 5 2 / 18
Hard 0 / 6 0 / 3 1 / 4 0 / 5 1 / 18
Gemini Flash Relevance 0 / 6 0 / 3 0 / 4 2 / 5 2 / 18
Diversity 0 / 6 0 / 3 0 / 4 2 / 5 2 / 18
Curriculum 0 / 6 0 / 3 0 / 4 0 / 5 0 / 18
Hard 0 / 6 0 / 3 0 / 4 0 / 5 0 / 18
Llama 3.1 Relevance 1 / 6 0 / 3 1 / 4 1 / 5 3 / 18
Diversity 0 / 6 0 / 3 0 / 4 2 / 5 2 / 18
Curriculum 0 / 6 0 / 3 0 / 4 1 / 5 1 / 18
Hard 0 / 6 0 / 3 0 / 4 2 / 5 2 / 18
Total Relevance 1 / 18 0 / 9 1 / 12 3 / 15 5 / 54
Diversity 0 / 18 0 / 9 1 / 12 6 / 15 7 / 54
Curriculum 1 / 18 0 / 9 0 / 12 2 / 15 3 / 54
Hard 0 / 18 0 / 9 1 / 12 2 / 15 3 / 54
Figure 3: Results with varying the number of examples for ICL with Gemini Pro, where we average the results for each task.
Results on Sample Selection Strategies
We report the detailed results of various sample selection approaches in many-shot ICL scenarios in Figure 2. To rigorously evaluate each sample selection approach and their statistically significant gains, we conduct a t-test with a 95% confidence threshold and report the results in Table 1. From these results, we observe that previously effective sample selection methods, designed for shorter context LMs, yield little to no performance gains over the random selection approach when applied to LCLMs. Aggregated results across three different LCLMs indicate statistical significance in fewer than 15% of instances, indicating that they are not reliable.
Analysis on Number of ICL Examples
To see the performance of ICL with respect to the number of examples, we visualize results in Figure 3. Overall, for any sampling method, we observe that performance increases as the number of examples increases. Also, when the number of examples is relatively small, the relevance-based sample selection approach performs particularly well, as focusing on highly relevant examples maximizes learning effectiveness when using a small number on examples. However, as the number of examples increases, the performance gap between various sample selection methods diminishes, indicating that performance is less dependent on selection strategies in many-shot scenarios. Lastly, in the summarization task (where samples tend to be longer than those in other tasks), we observe an initial increase in performance as more examples are added, followed by a decline once the context becomes heavily populated with a large number of examples. We argue this decline likely reflects the challenges LCLMs face in processing extremely long contexts, and we offer more analysis and discussion in Section 4.2.
Table 2: Results with varying the order of ICL samples, where Ascending and Descending represent cases where examples closer to the query appear earlier and later in the LCLM context, respectively. In contrast, random denotes the case where examples are arranged randomly without a specific order.
Methods Summarization Translation Reasoning Classification
Random 0.310 $\pm$ 0.004 0.553 $\pm$ 0.004 0.650 $\pm$ 0.023 0.539 $\pm$ 0.007
Ascending 0.307 $\pm$ 0.006 0.557 $\pm$ 0.004 0.641 $\pm$ 0.027 0.534 $\pm$ 0.010
Descending 0.309 $\pm$ 0.003 0.552 $\pm$ 0.007 0.648 $\pm$ 0.021 0.539 $\pm$ 0.005
Analysis on Example Order
Previous work has shown that earlier LMs are sensitive to the order of examples when doing few-shot ICL. For example, LMs tend to follow the answer in the last example Zhao et al. (2021); Lu et al. (2022). To investigate whether similar issues arise in many-shot ICL with LCLMs, we experiment by comparing performance when ordering ICL examples randomly, by increasing similarity, and by decreasing similarity. The results in Table 2 suggest that the order of examples does not affect performance of LCLMs.
Analysis on Computational Complexity
In addition to performance, computational complexity is a critical factor to consider when assessing the practicality of many-shot ICL with LCLMs, as they often handle million-token contexts. We note that for approaches that adjust ICL examples based on the given query (such as relevance-based selection), the complexity scales quadratically, $\mathcal{O}(n^{2})$, where $n$ represents the number of tokens used for ICL demonstrations. In contrast, the simpler naive selection approach, which uses the same set of randomly selected examples for all queries, offers a significantly more efficient complexity of $\mathcal{O}(kn)$, where $k$ is the number of tokens only within the target query ($n\gg k$). This is because the selected examples do not change based on the query; thus, the same set of examples can be key-value cached. As a result, random selection is a practical choice due to its equivalent performance with other selection methods and the added advantage of efficiency.
3 Augmenting ICL Demonstrations to Increase Context Capacity of LCLMs
Table 3: Results of LCLM-enabled ICL on four different tasks, where Random indicates the naive sample selection approach without selection criteria, Best Selection indicates the model that achieves the best performance among sophisticated sample selection methods for each experiment unit, and Augmentation indicates the proposed approach that generates demonstrations and uses them alongside original samples with random selection. We emphasize statistically significant results over Random in bold. We exclude Llama from the augmentation scenario as its context capacity is approximately ten times smaller than that of Gemini, allowing it to fully utilize its available context with the original examples alone, making augmentation unnecessary.
Translation Reasoning
LCLMs Methods ENG to BEM ENG to KMR ENG to EWE ENG to SPA ENG to FRA ENG to DEU Date Salient
Gemini Pro Random 0.470 $\pm$ 0.003 0.439 $\pm$ 0.001 0.419 $\pm$ 0.004 0.580 $\pm$ 0.006 0.734 $\pm$ 0.002 0.676 $\pm$ 0.010 0.854 $\pm$ 0.009 0.776 $\pm$ 0.035
Best Selection 0.470 $\pm$ 0.004 0.443 $\pm$ 0.004 0.418 $\pm$ 0.002 0.583 $\pm$ 0.004 0.745 $\pm$ 0.005 0.676 $\pm$ 0.004 0.896 $\pm$ 0.021 0.772 $\pm$ 0.017
[…]
Reasoning Classification All
LCLMs Methods Tracking7 Web Banking77 DialogRE Discovery FewNERD GoEmotion Average
Gemini Pro Random 0.294 $\pm$ 0.029 0.675 $\pm$ 0.021 0.878 $\pm$ 0.002 0.661 $\pm$ 0.009 0.195 $\pm$ 0.007 0.568 $\pm$ 0.012 0.393 $\pm$ 0.007 0.574 $\pm$ 0.010
Best Selection 0.311 $\pm$ 0.031 0.700 $\pm$ 0.028 0.886 $\pm$ 0.004 0.709 $\pm$ 0.014 0.204 $\pm$ 0.011 0.569 $\pm$ 0.006 0.413 $\pm$ 0.006 0.586 $\pm$ 0.011
[…]
Translation Reasoning
LCLMs Methods ENG to BEM ENG to KMR ENG to EWE ENG to SPA ENG to FRA ENG to DEU Date Salient
Gemini Flash Random 0.419 $\pm$ 0.006 0.427 $\pm$ 0.004 0.363 $\pm$ 0.002 0.573 $\pm$ 0.004 0.726 $\pm$ 0.004 0.666 $\pm$ 0.005 0.754 $\pm$ 0.022 0.682 $\pm$ 0.019
Best Selection 0.421 $\pm$ 0.002 0.434 $\pm$ 0.002 0.360 $\pm$ 0.003 0.575 $\pm$ 0.002 0.732 $\pm$ 0.003 0.673 $\pm$ 0.001 0.777 $\pm$ 0.030 0.687 $\pm$ 0.015
[…]
Reasoning Classification All
LCLMs Methods Tracking7 Web Banking77 DialogRE Discovery FewNERD GoEmotion Average
Gemini Flash Random 0.256 $\pm$ 0.030 0.582 $\pm$ 0.033 0.868 $\pm$ 0.004 0.541 $\pm$ 0.008 0.065 $\pm$ 0.007 0.521 $\pm$ 0.006 0.362 $\pm$ 0.016 0.520 $\pm$ 0.011
Best Selection 0.270 $\pm$ 0.031 0.566 $\pm$ 0.031 0.872 $\pm$ 0.006 0.547 $\pm$ 0.012 0.083 $\pm$ 0.007 0.532 $\pm$ 0.002 0.385 $\pm$ 0.006 0.528 $\pm$ 0.010
[…]
3.1 ICL Example Augmentation Approach
Recall that recent advances in LCLMs offer unprecedented context capacity, potentially amplifying ICL performance by including more examples. However, the available examples sometimes fall short of filling this expanded capacity, and this under-utilization of the context may result in suboptimal performance. To address this, we introduce a simple yet effective ICL sample augmentation approach designed to increase the context capacity of LCLMs, while being scalable for many-shot scenarios. This method consists of synthetic example generation and low-quality example filtering.
Generation of Synthetic Examples
Formally, let $\mathcal{D}=\{({\bm{x}}_{i},{\bm{y}}_{i})\}_{i=1}^{k}$ be a dataset of available ICL examples for a given target task, where each example $({\bm{x}}_{i},{\bm{y}}_{i})$ represents an input-output pair. The objective is to generate a set of synthetic examples $\mathcal{D}^{\prime}=\{({\bm{x}}^{\prime}_{j},{\bm{y}}^{\prime}_{j})\}_{j=1}^{m}$ (to supplement the original dataset $\mathcal{D}$), such that the augmented set of examples $\mathcal{D}_{\texttt{AUG}}=\mathcal{D}\cup\mathcal{D}^{\prime}$ can increase the utilization of the available context capacity of LCLMs. To operationalize this, we generate each synthetic example (${\bm{x}}^{\prime}_{j},{\bm{y}}^{\prime}_{j}$) by prompting an LM with randomly selected real examples from $\mathcal{D}$ as context, to ensure that the generated data retains meaningful patterns and characteristics relevant to the task.
Filtering Out Low-Quality Examples
Once the synthetic examples are generated, we filter out low-quality instances that may introduce noise or irrelevant information. To do this, we design a function $f$ that assigns a quality score to each synthetic example $({\bm{x}}^{\prime}_{j},{\bm{y}}^{\prime}_{j})$ based on its contextual relevance and alignment with real examples as well as overall quality. Specifically, each synthetic example is rated on a 5-point Likert scale by prompting the LM 30 times with the synthetic and 30 real examples. We then compute an aggregate score using a weighted average of scores with their corresponding probabilities from the LM. Only the synthetic examples that exceed a quality threshold, $\tau$, are retained in the augmented example set, as follows:
$\mathcal{D}_{\texttt{AUG}}=\mathcal{D}\cup\{({\bm{x}}^{\prime}_{j},{\bm{y}}^{\prime}_{j})\;|\;f({\bm{x}}^{\prime}_{j},{\bm{y}}^{\prime}_{j},\mathcal{D})\geq\tau\}_{j=1}^{m},$
where $f({\bm{x}}^{\prime}_{j},{\bm{y}}^{\prime}_{j},\mathcal{D})$ is the quality assessment function, and $\tau$ is the threshold value for filtering.
3.2 Experimental Setup
For synthetic data generation and filtering, we use Gemini Pro, one of the state-of-the-art LMs. We focus on tasks that underutilize the context capacity of LCLMs even when all available samples are provided, such as translation, reasoning, and classification. For each task, we generate 3,000 examples and retain only those with a quality score above the median among the generated samples. As a result, we use the original examples and 1,500 synthetic examples. The prompts used to elicit data generation and filtering are provided in Appendix A.
3.3 Experimental Results
Figure 4: Results with varying the ratio of noisy examples within the context of LCLMs, where we report the relative performance over the ICL without noisy examples (i.e., the noise ratio of 0) and the results are averaged over multiple runs.
Figure 5: Results across different percentages of context size utilized in LCLMs, where the x-axis represents the percentage of the full LCLM context used (according to the number of tokens over the full token length), and the y-axis shows the relative performance compared to the highest performance achieved for each dataset. Results are averaged over multiple runs.
Main Results
As shown in Table 3, which compares the example augmentation approach (with random selection) to other sample selection strategies, the augmentation approach demonstrates substantial performance gains across various datasets, which can be attributed to the greater diversity and volume of ICL examples achieved through synthetic data generation, leading to the effective utilization of the context capacity of LCLMs. Also, like the random selection approach, our augmentation method allows the reuse of the same examples across all queries. Thus, due to key-value caching, the augmentation approach is as efficient as random selection while achieving superior performance.
Table 4: Results on ablation study, where w/o Filtering and w/o Original denote the ICL results based on augmented samples without filtering and without original samples, respectively. Only Original is the performance without generated samples.
Methods Translation Reasoning Classification
Augmentation 0.571 $\pm$ 0.005 0.696 $\pm$ 0.027 0.560 $\pm$ 0.008
[…]
w/o Original 0.544 $\pm$ 0.002 0.611 $\pm$ 0.025 0.531 $\pm$ 0.007
[…]
Ablation Study on Augmentation
To see how each component in the augmentation approach contributes to performance gains, we conduct an ablation study. As shown in 3.3, we observe that the full augmentation method (called Augmentation), which uses both original examples and filtered synthetic examples in combination, achieves the best performance. In contrast, when the filtering step is omitted, performance decreases, indicating that filtering contributes positively by removing lower-quality synthetic examples. Also, a large performance drop occurs when original samples are excluded from the augmented set. This suggests that although filtering helps maintain quality, the synthetic samples generated still do not match the quality of the original examples. Thus, while our augmentation approach is effective, further research could improve data generation techniques to improve the quality of the synthetic examples.
4 Behaviors of LCLM-Enabled ICL
4.1 LCLM-Based ICL with Noisy Examples
LCLMs can accommodate a large number of diverse ICL examples, which raises the question of the impact and risk of including noisy examples in the context. We investigate how the performance of LCLM-enabled ICL is impacted when some or all of the ICL examples are noisy. To simulate noisy examples, we modify the outputs of a subset of in-context demonstrations by replacing their outputs with outputs from other randomly selected demonstrations. As shown in Figure 4, LCLM-enabled ICL is largely robust to noise when the proportion of noisy examples is relatively low (i.e., below 25%). This observation highlights why augmented examples, even if slightly lower quality, can still enhance performance as it increases the utilization of the context window. Also, when the amount of noise exceeds this threshold, LCLMs become vulnerable to the negative effects of noise and the performance notably declines. This adverse effect is more pronounced for challenging tasks, such as the GovReport dataset in summarization and low-resource translation tasks (e.g., English to Bemba or Ewe). This is likely because LCLMs are less familiar with those tasks, and therefore rely more on learning from in-content examples.
4.2 LCLM-Based ICL with Long Context
As the context length capacity of LCLMs continues to grow, it becomes increasingly important to assess whether LCLMs can reliably utilize a large number of ICL examples. To investigate this, we conduct an experiment analyzing the performance as a function of the context utilization. Specifically, we gradually increase the number of examples by powers of two, and if the entire set of examples within the dataset is used, we further extend the context utilization by repeating these examples. The hypothesis being tested is that if LCLMs can effectively understand and utilize extremely long context, performance should remain consistent even with repeated examples, as the presence of duplicates should not impact contextual understanding. However, as shown in Figure 5, a substantial performance decline occurs when LCLMs are pushed to use extremely large contexts. Specifically, this decline generally begins when more than 25% of the available context capacity is utilized. Also, the performance drop is pronounced in tasks such as xsum, which requires generating abstractive summaries (unlike other summarization datasets like arXiv or GovReport) and in tasks demanding complex reasoning such as date understanding (Date) and object tracking (Tracking7). These findings suggest that while LCLMs can handle moderately long contexts, they encounter limitations with exceedingly large contexts, particularly in tasks requiring fine-grained reasoning or abstractive generation. This may be due to challenges in distinguishing and integrating relevant information across numerous examples, especially when tasks require high levels of nuanced abstraction and precise reasoning.
5 Related Work
Long Context Language Models
The field of language modeling has witnessed remarkable advancements, particularly with the development of Large Language Models (LLMs) Brown et al. (2020); OpenAI (2023); Reid et al. (2024); Dubey et al. (2024). However, early LLMs were oftentimes constrained by relatively short context windows, typically handling only a few thousand tokens at a time, which limits their applicability in advanced tasks requiring broader context comprehension, such as document-level summarization or complex reasoning Koh et al. (2023); Suzgun et al. (2022). To address this, recent efforts have led to the development of Long Context Language Models (LCLMs), designed specifically to process much larger contexts, sometimes accommodating over a million tokens within a single prompt Reid et al. (2024). To mention a few, models like Longformer and BigBird Beltagy et al. (2020); Zaheer et al. (2020) incorporate sparse attention mechanisms to efficiently handle extended contexts without compromising on computational feasibility. Recent work has pushed these limits even further – for example, LongRoPE extends the the context window of LLMs to 2M tokens by interpolating their specific positional embeddings Ding et al. (2024).
In-Context Learning
In-Context Learning (ICL) is a recent paradigm that enables language models to learn from examples provided within their input context and then perform given tasks Brown et al. (2020); Min et al. (2022); von Oswald et al. (2023). Since its introduction, previous studies have concentrated on developing the strategies to optimize the quality and arrangement of in-context examples to maximize performance, especially given the limitations of early LMs on context length. For example, these approaches include selecting examples that maximize relevance to the target query Liu et al. (2021); Rubin et al. (2022), ensuring diversity among examples to cover a range of possible cases Sorensen et al. (2022); An et al. (2023), strategically ordering examples to improve model adaptation Zhao et al. (2021); Lu et al. (2022), and prioritizing examples by their ease of learning based on their difficultly Mavromatis et al. (2023); Liu et al. (2024). Yet, as the context capacity expands with LCLMs, these conventional selection strategies warrant re-evaluation, particularly in many-shot settings; thus, we focus on revisiting them.
Many-Shot ICL
Early approaches in many-shot ICL have primarily focused on the paradigm shift brought by the ability to incorporate a larger number of examples within the input context Agarwal et al. (2024); Bertsch et al. (2024), without giving much consideration to example selection strategies other than the random selection. Despite their simplicity, such many-shot ICL methods have sometimes demonstrated performance comparable to fine-tuning. Also, there is a very recent work that explores retrieval strategies in many-shot ICL Bertsch et al. (2024); however, they use models with relatively limited context capacities (e.g., under 100k tokens with Llama 2), resulting in restrictions on the number of examples included and, consequently, making retrieval-based methods appear more advantageous. However, contrary to this finding, we uncover that this advantage diminishes as the context capacity increases, allowing random sampling to perform on par with more sophisticated selection methods when a large number of examples is used. Lastly, other recent efforts include establishing benchmarks for long-context ICL Lee et al. (2024a); Li et al. (2024). Unlike prior studies, our work offers a novel perspective by systematically re-evaluating traditional selection strategies in the expanded context regime and highlighting the shift from selection optimization to effectively leveraging the extensive context space in many-shot ICL, with the proposal of data augmentation for cases where the number of examples is not sufficient to populate the context capacity of LCLMs.
6 Conclusion
We explored ICL in the context of LCLMs, which the enable inclusion of significantly more examples in-context than previously possible, and investigated whether traditional sample selection strategies remain effective in these many-shot scenarios. Through extensive experiments across diverse tasks and datasets, we observed that previously favored, sophisticated sample selection techniques offer minimal to zero performance gains over simple random selection in most cases. We believe this unexpected finding suggests a potential paradigm shift in ICL research: as LCLMs allow the processing of extensive contexts, sample selection may no longer be a priority, with simpler methods proving similarly effective and more computationally efficient due to key-value caching. We also highlighted the emerging challenge of underutilized context in low-resource tasks due to limited example availability, and, to address this, proposed a data augmentation strategy, which substantially boosts ICL performance by increasing context utilization of LCLMs. Lastly, we analyzed the behavior of LCLM-enabled ICL when operating with extremely long context and in the presence of noisy examples, and found that while performance improves with added examples, it plateaus and even declines when the context becomes too long, with increased vulnerability to noise in complex tasks. This suggests promising future directions in making LCLMs more robust to lengthy context and noise examples alongside the direction of extending their context length.
Limitations
While this work explores the new opportunity of ICL with LCLMs, a couple of limitations can be considered. First, the computational cost associated with LCLMs remains a significant challenge, particularly for researchers and practitioners in resource-constrained settings. Second, while the proposed data augmentation method enhances context utilization of LCLMs and improves ICL performance, the quality of synthetic examples often falls short of the quality of original data. Addressing them through cost-efficient strategies for leveraging LCLMs and developing improved data augmentation techniques would be an exciting area for future work.
Ethics Statement
We believe this work does not raise any direct ethical concerns, as it primarily focuses on advancing the understanding of ICL with LCLMs. However, as with any other application of LCLM-based ICL, careful consideration must be given to the quality of the examples used in the context. Specifically, the inclusion of biased, harmful, or otherwise problematic examples in the input context can propagate or amplify these issues in the model’s outputs, and we advise practitioners to carefully evaluate and select ICL examples to avoid potential issues.
Appendix A Prompts
We provide the prompts used for many-shot ICL on translation, summarization, and reasoning tasks in Table A and on classification tasks in Table A. Also, we provide the prompts used for synthetic data augmentation and filtering in Table A.
Table 5: A list of prompts that we use for many-shot ICL on translation, summarization, and reasoning tasks.
Types Prompts
Translation You are an expert translator. I am going to give you one or more example pairs of text snippets where the first is in {SOURCE_LANGUAGE} and the second is a translation of the first snippet into {TARGET_LANGUAGE}. The sentences will be written as the following format: {SOURCE_LANGUAGE}: <first sentence> {TARGET_LANGUAGE}: <translated first sentence> After the example pairs, I am going to provide another sentence in {SOURCE_LANGUAGE} and I want you to translate it into {TARGET_LANGUAGE}. Give only the translation, and no extra commentary, formatting, or chattiness. Translate the text from {SOURCE_LANGUAGE} to {TARGET_LANGUAGE}. {EXAMPLES} {TARGET_QUERY}
[…]
Table 6: A list of prompts that we use for many-shot ICL on five different extreme classification tasks.
Types Prompts
BANKING77 I am going to give you one or more example pairs of customer service query and its intent. The pairs will be written as the following format: service query: <query> intent category: <category> After the example pairs, I am going to provide another customer service query and I want you to classify the label of it that must be one among the intent categories provided in the examples. Give only the category, and no extra commentary, formatting, or chattiness. {EXAMPLES} {TARGET_QUERY}
[…]
Table 7: A list of prompts that we use for generating synthetic demonstrations and filtering them of low-quality.
Types Prompts
Generation You are an expert in data augmentation. You will be provided with a series of demonstrations that show how a task is performed. Your objective is to generate a new example that closely follows the pattern, structure, and style of the demonstrations. Carefully analyze the key steps, transitions, and output style in the provided demonstrations. Then, create a new sample that maintains consistency in format and correctness while introducing variety in content. Here are the demonstrations: {EXAMPLES} Now, as an expert, generate a new sample that aligns with the original demonstrations:
[…]